Northwoods Backpackers is a retail catalog store in Vermont that specializes in outdoor clothing and camping equipment. Phone orders are taken each day by a large pool of computer operators, some of whom are permanent and some temporary. A permanent operator can process an average of 76 orders per day, whereas a temporary operator can process an average of 53 orders per day. The company averages at least 600 orders per day. The store has 10 computer workstations. A permanent operator processes about 1.3 orders with errors each day, whereas a temporary operator averages 4.1 orders with errors daily. The store wants to limit errors to 24 per day. A permanent operator is paid $81 per day, including benefits, and a temporary operator is paid $50 per day. The company wants to know the number of permanent and temporary operators to hire to minimize costs.
Answer #1

Let P and T represent the number of permanent and temporary workers to be employed.

Minimize Z = total cost = 81 P + 50 T

Subject to,

76 P + 53 T >= 600 (minimum orders to be processed)

1.3 P + 4.1 T <= 24 (maximum number of errors)

P + T <= 10 (maximum number of workstations)

P, T >= 0, Integers

Excel formulation

F G Parmanent Temporary 2 workers workers No. of Daily cost 4 81 50 SUMPRODUCT(SCS3:SD$3,C4.D4) Min total orders 76 Max error

Set Objective OMin Value Of Max By Changing Variable Cells: $C$3:$D$3 Subject to the Constraints: $C$3:SD$3-integer Add Chang

Solution

Objective Cell (Min) Original Vae Final Value 0 Cell Name SE$A Daily cost 636 Variable Cells Original Va Name No. of Parmanen